### v1

# When Knockoffs fail: diagnosing and fixing non-exchangeability of Knockoffs

Abstract
Knockoffs are a popular statistical framework that addresses the challenging problem of conditional variable selection in high-dimensional settings with statistical control. Such statistical control is essential for the reliability of inference. However, knockoff guarantees rely on an exchangeability assumption that is difficult to test in practice, and there is little discussion in the literature on how to deal with unfulfilled hypotheses. This assumption is related to the ability to generate data similar to the observed data. To maintain reliable inference, we introduce a diagnostic tool based on Classifier Two-Sample Tests. Using simulations and real data, we show that violations of this assumption occur in common settings for classical Knockoffs generators, especially when the data have a strong dependence structure. We show that the diagnostic tool correctly detects such behavior. To fix knockoff generation, we propose a nonparametric, computationally-efficient alternative knockoff construction, which is based on constructing a predictor of each variable based on all others. We show that this approach achieves asymptotic exchangeability with the original variables under standard assumptions on the predictive model. We show empirically that the proposed approach restores error control on simulated data.
Controlled variable selection,
Knockoffs,
Classifier Two-Sample tests,
keywords:
[…]
, , and
1 Introduction
Controlled variable selection is a fundamental problem encountered in diverse fields where practitioners have to assess the importance of input variables with regards to an outcome of interest. Variable selection aims at choosing variables that should be included in an analysis or model while excluding those that do not substantially contribute to the understanding or prediction of an outcome. Statistically controlled variable selection aims at limiting the proportion of selected variables that are independent of the outcome. This problem is much harder than usual variable selection since guarantees on False Positives have to be provided.
For instance, selecting variables is useful in epidemiology to better understand driving factors in the propagation of diseases (Greenland, 1989). In mathematical finance, choosing an optimal portfolio amounts to solving a variable selection problem (Aït-sahali and Brandt, 2001). In neuroimaging, practitioners are interested in associating cognitive functions to relevant brain regions (Weichwald et al., 2015). In genomics, identifying gene variants that drive disease outcomes is crucial (Visscher et al., 2012). Additionally, False Discoveries have to be controlled to ensure that variable selection is reliable (Benjamini and Hochberg, 1995; Genovese and Wasserman, 2004).
In conditional variable selection, the objective is to assess the value of a particular variable w.r.t. an outcome given a certain set of other variables (König et al., 2021). This approach differs markedly from classical marginal inference, where variables are considered individually in relation to the outcome (Candès et al., 2018). With conditional variable selection, we examine whether a variable remains informative and relevant when considered alongside other variables. This is particularly crucial when dealing with datasets characterized by high levels of correlation among variables. Inference in the presence of correlations is illustrated with a genomics example of Genome-Wide Association Study (GWAS) in Figure 1, where correlation between variables is due to a biological phenomenon called linkage disequilibrium (Uffelmann et al., 2021).
Figure 1: Illustration of conditional variable selection. Given a genomic dataset and kidney stones diagnoses (outcome), we aim at finding variables that add information w.r.t the outcome given all other variables. This problem is challenging since genes are highly correlated locally. Assessing the added information of specific genes given all others is crucial to improve the understanding of disease outcomes. Details about this Manhattan plot are available in Howles et al., 2019.
This problem becomes difficult when there is a high level of correlation between variables on which one aims to perform inference. In machine learning contexts, the difficulty of identifying features that add unique predictive value when combined with others in large-correlation contexts has been well identified (Bickel, Ritov and Tsybakov, 2009; Mandozzi and Bühlmann, 2016; Goeman and Bühlmann, 2007; Javanmard and Montanari, 2014). High correlations are ubiquitous in many application areas, for example when the features of interest are derived from some biological measurements that inherently have complex dependencies. For instance, in brain mapping, practitioners may want to detect brain regions that are relevant for a given cognitive task given the rest of the brain (Weichwald et al., 2015), but have to deal with the strong dependencies observed in these data (Chevalier et al., 2021). In genomics, understanding which genes conditionally affect disease outcomes can lead to more effective disease prevention and intervention strategies (Sesia, Sabatti and Candès, 2019).
Knockoffs (Barber and Candès, 2015; Candès et al., 2018) have emerged as a popular and powerful tool for addressing the problem of conditional variable selection with guarantees on the rate of false detections. This framework consists of guiding variable selection by generating synthetic variables, referred to as knockoffs, that closely mimic the statistical properties of the original variables while not being associated to the outcome of interest (conditionally on their original counterpart). This approach aims to identify genuine associations between variables and outcomes by comparing them with their corresponding knockoffs. An attractive feature of knockoffs is that inference is performed simultaneously for all variables, resulting in an acceptable computational cost. Performing the same type of inference for each variable leads to as many Conditional Randomization Tests (Candès et al., 2018; Nguyen, Thirion and Arlot, 2022) as the number of covariates, which has a high computational cost.
It is important to note, however, that the error control of knockoffs relies on a critical assumption called exchangeability. For exchangeability to hold, the joint distribution of the data must remain unchanged when an original variable is exchanged for its knockoff counterpart. While Knockoffs have shown promise in many applications, this assumption of exchangeability requires careful consideration and assessment, as its validity impacts the reliability of the entire variable selection process. While the generative processes could ideally be known or derived from first principles, for many practical problems, they are unknown or intractable ; current practice then relies on learning the joint distributions of the observations in order to draw knockoffs from an appropriately perturbed model.
The most popular knockoff generation procedure involves a Gaussian assumption, and thus relies on the knowledge or the accurate estimation of the covariance structure of the covariates, which leads to two potential issues: i) the violation of the Gaussian hypotheses and ii) inaccuracies in the covariance estimation. This worsens with the number of variables $p$ as the problem becomes harder, especially in high-dimensional regimes where the number of samples $n$ is comparatively small. Some alternative Knockoff generation methods avoid the problem of covariance estimation via e.g. Deep Learning (Romano, Sesia and Candès, 2020). However, such methods also suffer from high-dimensional regimes as deep neural networks require massive amounts of data to be properly trained, especially in large feature spaces. Overall, methods that require a large sample size $n$ and a small number of variables $p$ may not be adapted to the Knockoffs framework, which is designed for high-dimensional variable selection. Domain-specific procedures have been developed to tackle some of these issues. For instance, in the field of genomics where variables of interest may be discrete, Knockoffs can be built using Hidden Markov Models (HMM) (Rabiner and Juang, 1986), as demonstrated in the work of Sesia, Sabatti and Candès (2019).
The aim of this work is twofold: First, we propose an effective diagnostic tool that allows practitioners to examine knockoffs along the original data for potential violations. Based on this, we highlight common cases of non-exchangeability when using Gaussian knockoffs and associated error control violations. We describe their consequences, typically a failure to control the false discovery rate at the nominal level. We propose an efficient non-parametric algorithm for constructing knockoffs which comes with theoretical guarantees. We show through simulations that this procedure restores error control in all cases considered.
2 Knockoffs framework and related work
Notation. For an integer $k$, $\llbracket k\rrbracket$ denotes the set $\{1,\dots,k\}$. Equality in distribution is denoted by =$\stackrel{{\scriptstyle d}}{{=}}$. We denote vectors by bold lowercase letters. A vector $\mathbf{x}=(x_{j})_{j\in\llbracket p\rrbracket}$ from which the $j^{th}$ coordinate is removed is denoted by $\mathbf{x}_{-j}$, i.e. $\mathbf{x}\setminus\{x_{j}\}$. Independence between two random vectors $\mathbf{x}$ and $\mathbf{y}$ is denoted by $\mathbf{x}\perp\mathbf{y}$. For two vectors $\mathbf{x}$ and $\mathbf{\tilde{x}}$ of size $p$ and a subset $S\subset\llbracket p\rrbracket$, $(\mathbf{x},\mathbf{\tilde{x}})_{swap(S)}$ denotes the vector obtained from $(\mathbf{x},\mathbf{\tilde{x}})$ by swapping the entries $x_{j}$ and $\tilde{x}_{j}$ for each $j\in S$. Matrices are denoted by bold uppercase letters, the only exception being the vector of Knockoff statistics that we denote by $\mathbf{W}$ as in Barber and Candès (2015); Candès et al. (2018). For any set $S$, $|S|$ denotes the cardinality of $S$. For a vector $\mathbf{z}=(z_{j})_{1\leq j\leq p}$ and $S\subset\llbracket p\rrbracket$, we denote by $z_{(j:S)}$ (or $z_{(j)}$ when there is no ambiguity) the $j^{th}$ smallest value in the sub-vector $(\mathbf{z}_{s})_{s\in S}$.
Problem setup. The input data are denoted by $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n$ is the number of samples and $p$ the number of variables. The outcome of interest is denoted by $\mathbf{y}\in\mathbb{R}^{n}$. The goal is to select variables that are relevant with regards to the outcome conditionally on all others. Formally, we test simultaneously for all $j\in\llbracket p\rrbracket$:
$H_{0,j}:y\perp x_{j}|\mathbf{x}_{-j}\quad\text{versus}\quad H_{1,j}:y\not\perp x_{j}|\mathbf{x}_{-j}.$
The output of a variable selection method is a rejection set $\hat{S}\subset\llbracket p\rrbracket$ that estimates the true unknown support $\mathcal{H}_{1}=\{j:y\not\perp x_{j}|\mathbf{x}_{-j}\}$. Its complement is the set of true null hypotheses $\mathcal{H}_{0}=\{j:y\perp x_{j}|\mathbf{x}_{-j}\}$. To ensure reliable inference, our aim is to provide a statistical guarantee on the proportion of False Discoveries in $\hat{S}$. The False Discovery Proportion (FDP) and the False Discovery Rate (FDR) Benjamini and Hochberg (1995) are defined as:
$\mathrm{FDP}(\hat{S})=\frac{|\hat{S}\cap\mathcal{H}_{0}|}{|\hat{S}|\vee 1},\quad\mathrm{FDR}(\hat{S})=\mathbb{E}[\mathrm{FDP}(\hat{S})].$
Knockoffs. Knockoff inference leverages noisy duplicates known as knockoff variables which serve as controls in the variable selection process. A key challenge in this method is to ensure that these knockoff variables maintain the same correlation structure as the original variables, while being conditionally independent of the outcome. This is essential to enable meaningful comparisons between the original variables and their knockoff counterparts, thereby identifying variables that provide relevant information regarding the outcome. Knockoff variables are defined as follows:
Definition 1 (Model-X Knockoffs, Candès et al., 2018).
For the family of random variables $\mathbf{x}=(x_{1},\ldots,x_{p})$, Knockoffs are a new family of random variables $\mathbf{\tilde{x}}=(\tilde{x}_{1},\ldots,\tilde{x}_{p})$ satisfying:
  1. 1.
    for any $S\subset\llbracket p\rrbracket$, $(\mathbf{x},\mathbf{\tilde{x}})_{swap(S)}\stackrel{{\scriptstyle d}}{{=}}(\mathbf{x},\mathbf{\tilde{x}})$
  2. 2.
    $\mathbf{\tilde{x}}\perp\mathbf{y}|\mathbf{x}$.
Exchangeability is the first property of this definition. For any set of swapped variables, the joint distribution of real and knockoff variables must remain identical to the original one. Building Knockoff variables that satisfy this property is a difficult problem in itself – we discuss existing work on Knockoff generation and propose a novel approach in Section 3.
Statistical inference on Knockoff variables. Once valid knockoffs have been created, conditional variable selection can be performed. To distinguish variables that are substantially more important than their corresponding Knockoffs, a machine learning model is employed to generate importance scores for each variable and its respective knockoff. This step enables the identification of variables that offer valuable insights into the outcome, as they exhibit significant disparities in importance compared to their knockoffs. Quantitatively, this is done by computing Knockoff statistics $\mathbf{W}=(W_{1},\ldots,W_{p})$ that are defined as follows.
Definition 2 (Knockoff Statistic, Candès et al., 2018).
A knockoff statistic $\mathbf{W}=(W_{1},\ldots,W_{p})$ is a measure of feature importance that satisfies:
  1. 1.
    $\mathbf{W}$ depends only on $\mathbf{X},\mathbf{\tilde{X}}$ and $\mathbf{y}$: $\mathbf{W}=g(\mathbf{X},\mathbf{\tilde{X}},\mathbf{y})$.
  2. 2.
    Swapping column $\mathbf{x}_{j}$ and its knockoff column $\mathbf{\tilde{x}}_{j}$ switches the sign of $W_{j}$:
    $W_{j}([\mathbf{X},\mathbf{\tilde{X}}]_{swap(S)},\mathbf{y})=\left\{\begin{array}[]{ll}W_{j}([\mathbf{X},\mathbf{\tilde{X}}],\mathbf{y})\ \text{if }j\in S^{c}\\ -W_{j}([\mathbf{X},\mathbf{\tilde{X}}],\mathbf{y})\ \text{if }j\in S.\\ \end{array}\right.$
The most commonly used Knockoff statistic is the Lasso-coefficient difference (LCD) (Weinstein et al., 2020). This statistic is obtained by fitting a Lasso estimator (Tibshirani, 1996) on $[\mathbf{X},\mathbf{\tilde{X}}]\in\mathbb{R}^{n\times 2p}$, which yields $\widehat{\boldsymbol{\beta}}\in\mathbb{R}^{2p}$. Then, the Knockoff statistic can be computed using $\widehat{\boldsymbol{\beta}}$:
$\forall j\in\llbracket p\rrbracket,\quad W_{j}=|\widehat{\beta}_{j}|-|\widehat{\beta}_{j+p}|.$
This coefficient summarizes the importance of the original $j^{th}$ variable relative to its own Knockoff: $W_{j}>0$ indicates that the original variable is more important for fitting $y$ than the Knockoff variable, meaning that the $j^{th}$ variable is likely relevant. Conversely, $W_{j}<0$ indicates that the $j^{th}$ variable is probably irrelevant. We thus wish to select variables corresponding to large and positive $W_{j}$. Formally, the rejection set $\hat{S}$ of the Knockoffs method can be written $\hat{S}=\{j:W_{j}>T_{q}\},$ where:
$T_{q}=\min\left\{t>0:\frac{1+\#\left\{j:W_{j}\leqslant-t\right\}}{\#\left\{j:W_{j}\geqslant t\right\}}\leqslant q\right\}.$
This definition of $T_{q}$ ensures that the $\operatorname{FDR}$ is controlled at level $q$ (Candès et al., 2018). Alternatively, inference can be performed using $\pi$-statistics, which quantify the evidence against each variable:
$\pi_{j}=\left\{\begin{array}[]{l}\frac{1+|\left\{k:W_{k}\leq-W_{j}\right\}|}{p}\text{ if }\quad W_{j}>0\\ 1\quad\text{ if }\quad W_{j}\leq 0.\end{array}\right.$ (1)
As noted by Nguyen et al. (2020), the vanilla Knockoffs procedure of Candès et al., 2018 is equivalent to using the Benjamini and Hochberg (1995) procedure at level $q$ on the vector of $\pi$-statistics $(\pi_{j})_{j\in\llbracket p\rrbracket}$. The complete procedure is summarized in Figure 2.
Figure 2: Conditional variable selection via Knockoff variables. Knockoff variables are noisy copies of the original variables that are used as controls for variable selection (Panel a). Conditionally on the real variables, knockoffs are independent of the outcome $y$. Importance scores for each variable are computed using the sensitivity estimate from machine learning model – typically, the corresponding coefficient of a Lasso fit. The Knockoff statistic is defined as the difference of modulus between the score associated to the real variable and its Knockoff counterpart (Panel b). Then, $\pi$-statistics are computed to rank variables and to perform inference (Panel c).
Randomness. Beyond the validity of Knockoff variables, another notable challenge is the random nature of the knockoffs generation process. This randomness introduces instability in variable selection outcomes – this is highly undesirable in applications and hampers reproducibility.
Much work has been done recently to tackle this issue, including using stability selection criteria (Ren, Wei and Candès, 2021), aggregating multiple $\pi$-statistics (Nguyen et al., 2020; Blain et al., 2023) and using aggregated Knockoffs e-values (Ren and Barber, 2022). The gist of these approaches is to draw multiple sets of Knockoffs and aggregate the resulting statistics or rejection sets in order to robustify the variable selection process. Such methods offer ways to provide more stable and reliable selection sets, thus mitigating the randomness associated with the knockoff generation process. These advancements enhance the credibility and reproducibility of conditional variable selection using knockoffs.
3 Knockoffs generation methods
3.1 Gaussian Knockoffs
Candès et al., 2018 have built Knockoffs that are provably exchangeable for Gaussian data. Provided an observed $X$ with $X\sim\mathcal{N}(0,\Sigma)$, we sample from
$\tilde{X}\mid X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\mathcal{N}(\mu,\mathbf{V}),$
where
$\begin{gathered}\mu=X-X\boldsymbol{\Sigma}^{-1}\operatorname{diag}(s)\\ \mathbf{V}=2\operatorname{diag}(s)-\operatorname{diag}(s)\boldsymbol{\Sigma}^{-1}\operatorname{diag}(s)\end{gathered}$
With $\operatorname{diag}(s)$ any diagonal matrix. Then, if $\operatorname{diag}(s)$ is chosen such that $\mathbf{G}$ is positive semidefinite, we obtain valid Knockoffs since:
$[X,\tilde{X}]\sim\mathcal{N}(0,\mathbf{G}),\quad\text{ where }\mathbf{G}=\left(\begin{array}[]{cc}\boldsymbol{\Sigma}&\boldsymbol{\Sigma}-\operatorname{diag}(s)\\ \boldsymbol{\Sigma}-\operatorname{diag}(s)&\boldsymbol{\Sigma}\end{array}\right)$
A major challenge is that the covariance matrix $\boldsymbol{\Sigma}$ is unknown and has to be estimated via shrinkage procedures such as Ledoit-Wolf shrinkage (Ledoit and Wolf, 2003) or Graphical Lasso (Friedman, Hastie and Tibshirani, 2008). Given a correctly estimated $\boldsymbol{\Sigma}$, examples of valid $s$ include the equicorrelated construction $s_{j}^{\mathrm{EQ}}=2\lambda_{\min}(\Sigma)\wedge 1$ for all $j$.
In practice, testing exchangeability for non-Gaussian data is challenging. In the case of Gaussian data, since the actual covariance is not known but has to be estimated, exchangeability depends directly on the accuracy of the covariance estimate. Even if the data is Gaussian, its covariance matrix may be hard to estimate properly in high-dimensional regimes (Stein, Efron and Morris, 1972; Fukunaga, 2013). Indeed, such covariance estimation has to draw a difficult compromise between data fit and positive definiteness, along sparsity of the inverse covariance estimation from the Graphical Lasso. L2 shrinkage covariance estimation (Ledoit and Wolf, 2003) on the other hand, is known to lead to excessive bias (Belilovsky et al., 2017). In addition to the sheer computation cost, hyperparameter setting for the Graphical Lasso is challenging, and leads to difficult and costly parameter selection.
3.2 Existing alternatives to Gaussian Knockoffs
Other available methods to build Knockoffs include Monte-Carlo based methods (Sesia, Sabatti and Candès, 2019; Bates et al., 2021). These procedures focus on discrete distributions as in the case of GWAS studies. In the continuous setting, the method proposed in Bates et al., 2021 is equivalent to Gaussian Knockoffs.
Recent machine learning techniques such as deep learning have also been used to apprehend complex data dependence structures in the context of Knockoffs. Romano, Sesia and Candès, 2020 propose using Deep Neural networks to build Knockoffs using a target covariance matrix estimated from the data as in the Gaussian algorithm. Additionally, the proposed algorithm aims at minimizing the correlation between original variables and Knockoffs to maximize power. Using similar ideas, Zhu et al., 2021; Liu and Zheng, 2018 propose using a Variational Auto-Encoders (VAE) to build Knockoffs as VAEs allow learning and sampling from complex data distributions. However, in the context of high-dimensional variable selection, the number of samples available may be insufficient to properly train a Deep Learning model. A limitation of these approaches is that they offer no theoretical guarantees on Knockoffs exchangeability.
3.3 An efficient Non-parametric Knockoff generation algorithm
We propose an alternative method to build Knockoffs by learning to predict $X_{j}$ from $X_{-j}$. This idea is related to the Sequential Conditional Independent Pairs algorithm of Candès et al. (2018). A related idea is exploited to quantify variable importance in Random Forests in Chamma, Engemann and Thirion, 2023. The proposed approach is described in Algorithm 1.
In this approach, the dependency structure of the data is learned using a machine learning model $f$ for each variable that predicts $X_{j}$ from $X_{-j}$. Typically, using a Lasso model is a good default choice. In practice, we set $\lambda=\lambda_{max}/100$ with $\lambda_{max}=\frac{1}{n}\left\|\mathbf{X_{-j}}^{T}X_{j}\right\|_{\infty}$. The $j$-th fitted predictor is denoted $f_{j}$. Then, the residual $\widehat{\epsilon}_{j}=X_{j}-f_{j}(\mathbf{X_{-j}})$ is computed. Once this is done for all variables, Knockoffs are built by shuffling the residuals: $\tilde{X}_{j}\text{ is chosen as }f_{j}(\mathbf{X_{-j}})+\widehat{\epsilon}_{\tau{(j)}}$ with $\tau$ a permutation of $\llbracket p\rrbracket$. The important point is that this approach circumvents the difficult problem of covariance estimation in high-dimensional settings. Note that using a Lasso model is suitable in regimes where $n\simeq p$ – if $n\gg p$ selecting a more expressive model such as random forests (Breiman, 2001) or gradient boosting trees (Chen et al., 2015) is necessary to retain control.
Input : $f$
1 for $j\in[1,p]$ do
2      Fit a prediction model $f_{j}$ on $(\mathbf{X_{-j}},X_{j})$ // Typically a Lasso model
3       Compute the residual $\widehat{\epsilon}_{j}=X_{j}-f_{j}(\mathbf{X_{-j}})$
4 end for
5
6for $j\in[1,p]$ do
7      Sample $\text{Sample }\tilde{X}_{j}=f_{j}(\mathbf{X_{-j}})+\widehat{\epsilon}_{\tau{(j)}}$ // $\tau$ is a permutation of $\llbracket p\rrbracket$
8
9 end for
Return $\mathbf{\tilde{X}_{1:p}}$
Algorithm 1 Generating Non-parametric Knockoffs by learning to predict $X_{j}$ from $\mathbf{X_{-j}}$ using a model $f$.
We now turn to theoretically grounding the proposed approach. To tackle this problem, we use a metric formulation of the exchangeability condition defined in Definition 1.
Remark 1.
Let $\mathit{d}^{*}$ a distance over distributions of $\mathbb{R}^{2p}$ and $S\subset\llbracket p\rrbracket$. Then, exchangeability can be written using $\mathit{d}^{*}$:
[…]
We can now state the Knockoffs validity result, which is proven in Appendix A:
Theorem 1.
Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ the original data and $\mathbf{\tilde{X}}\in\mathbb{R}^{n\times p}$ Knockoffs built using Algorithm 1 with prediction model $f$. Assume that $\forall j\in\llbracket p\rrbracket$, the prediction error of $f_{j}$ converges to 00, i.e. that $\forall j\in\llbracket p\rrbracket,\left\|\widehat{\boldsymbol{\epsilon_{j}}}-\boldsymbol{\epsilon_{j}}\right\|_{2}^{2}\xrightarrow[n\to\infty]{}0$. Then, Knockoffs and original variables are asymptotically exchangeable in the number of samples $n$:
$\forall S\subset\llbracket p\rrbracket,\mathit{d}^{*}((\mathbf{X},\mathbf{\tilde{X}})_{swap(S)},(\mathbf{X},\mathbf{\tilde{X}}))\xrightarrow[\enskip n\to\infty\enskip]{}0$
Remark 2.
When using a Lasso model for $f$, the prediction error converges to 00 under standard assumptions described in Theorem 11.2 of Hastie, Tibshirani and Wainwright, 2015. In turn, Theorem 1 holds when using a Lasso model.
4 When Knockoffs fail: diagnosing non-exchangeability
A direct consequence of non-exchangeability is that Knockoff statistics of null variables ${\{X_{j}\text{ for }j\in H_{0}\}}$ are no longer symmetrical – yet this property is key to achieving error control via valid Knockoffs as in Candès et al., 2018. Said otherwise, non-exchangeability can make Knockoff importance scores non comparable with real variable importance scores which leads to bias in Knockoff statistics. This is illustrated in Figure 3 using a simulated data setup described in Section 5.
On the left panel, we display the inverse Cumulative Distribution Function (CDF) of Knockoff statistics of null variables, where Knockoffs come from an oracle, with known covariance. On the central panel, we display the inverse CDF in the same setup but with a covariance estimate that relies on the Graphical Lasso. On the right panel, we display the inverse CDF of statistics obtained using non-parametric Knockoffs built from the data. Knockoff statistics of oracle Knockoffs are nearly perfectly symmetric, as seen on the left panel. Notice that Knockoff statistics of null variables using data-derived Gaussian Knockoffs are not symmetric: they are skewed towards positive values, signalling that real and Knockoff importance scores are not comparable. Non-parametric Knockoffs nearly recover exact symmetry of null Knockoffs statistics.
Figure 3: Inverse CDFs of Knockoff statistics of null variables for oracle and data-derived Knockoffs. On the left panel, we display Knockoff statistics of null variables in an oracle simulation setup where perfectly valid Knockoffs have been built – i.e., using the true covariance of the data, computed using a massive amount of samples. Notice that the inverse CDF is practically symmetrical, which ensures validity of the inference. On the central panel, we display Knockoff statistics of null variables in the same setup but with a covariance estimate that relies on the Graphical Lasso. Notice that null Knockoffs statistics are not symmetrical: they are skewed towards positive values, signalling that real and Knockoff importance scores are not comparable. On the right panel, we display Knockoff statistics of null variables using non-parametric Knockoffs. Notice that symmetry is recovered, which ensures reliable inference. We use $n=500$ samples and $p=500$ variables.
In the remainder of this section, we will focus on diagnosing this problem before assessing its consequences on error control in practice.
A sufficient condition for non-exchangeability. Clearly, Knockoffs exchangeability does not hold if $\mathbf{x}$ is non-Gaussian while $\mathbf{\tilde{x}}$ is Gaussian. One can simply take $S=\llbracket p\rrbracket$; then $(\mathbf{x},\mathbf{\tilde{x}})_{swap(S)}=(\mathbf{\tilde{x}},\mathbf{x})\stackrel{{\scriptstyle d}}{{\neq}}(\mathbf{x},\mathbf{\tilde{x}})$. More broadly, any Knockoff generation procedure that relies on unfulfilled assumptions on variable distribution or dependence structure fails to replicate the distribution of $\mathbf{x}$. This leads to a natural sufficient condition to diagnose non-exchangeability: if there exists a classifier that is able to accurately distinguish samples from $\mathbf{x}$ versus samples from $\mathbf{\tilde{x}}$, then $\mathbf{\tilde{x}}\neq\mathbf{{x}}$ in distribution and exchangeability is violated. This idea is related to the C2ST (Classifier Two-Sample Testing) literature (Gretton et al., 2012; Lopez-Paz and Oquab, 2016).
Classifier Two-Sample Testing for Knockoffs. Formally, we wish to test the null hypothesis $H_{0}:\mathbf{\tilde{x}}\stackrel{{\scriptstyle d}}{{=}}\mathbf{{x}}$ given $n$ samples of each distribution. Following Section 3 of Lopez-Paz and Oquab, 2016, we proceed by constructing the dataset
$\mathcal{D}=\left\{\left(x^{i},0\right)\right\}_{i=1}^{n}\cup\left\{\left(\tilde{x}^{i},1\right)\right\}_{i=1}^{n}=:\left\{\left(z^{i},l^{i}\right)\right\}_{i=1}^{2n}.$
Then, the $2n$ samples of $\mathcal{D}$ are shuffled at random and and split into disjoint training and testing subsets $\mathcal{D}_{\text{tr }}$ and $\mathcal{D}_{\text{te }}$, where $\mathcal{D}=\mathcal{D}_{\text{tr }}\cup\mathcal{D}_{\text{te }}$ and $n_{\text{te }}:=\left|\mathcal{D}_{\text{te }}\right|$. Note that in practice, this split is performed several times to mitigate randomness, as in classical cross-validation.
Definition 3 (C2ST statistic, Lopez-Paz and Oquab, 2016).
Given $\mathcal{D}_{\text{tr }}$ and $\mathcal{D}_{\text{te }}$, where $\mathcal{D}=\mathcal{D}_{\text{tr }}\cup\mathcal{D}_{\text{te }}$ with $n_{\text{te }}:=\left|\mathcal{D}_{\text{te }}\right|$ and a binary classifier $g:\mathcal{R}^{p}\rightarrow\{0,1\}$ trained on $\mathcal{D}_{\text{tr}}$ the C2ST statistic $\hat{t}$ is defined as the classification accuracy on $\mathcal{D}_{\text{te }}$:
$\hat{t}=\frac{1}{n_{\text{te }}}\sum_{\left(z^{i},l^{i}\right)\in\mathcal{D}_{\text{te }}}1\left[g\left(z^{i}\right)=l^{i}\right]$
Intuitively, if $\mathbf{\tilde{x}}\stackrel{{\scriptstyle d}}{{=}}\mathbf{{x}}$, the test accuracy $\hat{t}$ should remain near 0.5, corresponding to chance-level. Conversely, if $\mathbf{\tilde{x}}\stackrel{{\scriptstyle d}}{{\neq}}\mathbf{{x}}$ and that the binary classifier is able to unveil distributional differences between the two samples, the test classification accuracy $\hat{t}$ should be greater than chance-level. The procedure is summarized in Figure 4.
Figure 4: Using a Classifier Two-Sample Test to diagnose exchangeability problems in Knockoffs. We input both input data and Knockoffs to a binary classifier with two different labels – the goal is to check whether a classifier can distinguish between real variables and Knockoff variables. If the classifier’s accuracy is substantially above chance level, exchangeability is violated since the joint distributions of input data and Knockoffs are not equal.
Exchangeability violation by improper sample pairing. Note that the C2ST diagnostic tool only tests that Knockoffs are indeed sampled from the same distribution as the original data. In practice, to have valid Knockoffs, a consistent pairing between original and Knockoffs samples is also needed. Said otherwise, knowing how to sample from the original data distribution is not enough to build valid Knockoffs. We illustrate this point via a simple experiment in the same simulation setup described in Section 5. The experiment consists in shuffling $50\%$ of the sample pairings of valid Knockoffs. The middle panel reports the FDP for each simulation run in both scenarios (shuffled and non-shuffled). The left and right sketches schematically illustrate shuffling sample pairings in $2D$. Note that the C2ST diagnostic is invariant to this shuffling.
Figure 5: Exchangeability violation by improper sample pairing. In the left boxplot, we build valid Knockoffs via the Gaussian algorithm and perform inference as in (Candès et al., 2018). In the right boxplot, we use the same Knockoffs but shuffle $50\%$ of the sample pairings before inference. We repeat this experiment $10$ times. Note that, coherently with theory, the $FDR$ is indeed controlled in the left boxplot. However, in the right boxplot, error control is lost due to improper sample pairings. The left and right sketches schematically illustrate shuffling sample pairings in $2D$.
As seen in Figure 5, FDR control is lost when shuffling $50\%$ of the sample pairings of valid Knockoffs. This shows that, beyond equality of the distributions, a proper sample pairing is needed to obtain valid Knockoffs. To test this in practice, one may compute the optimal assignment between real samples and Knockoff samples e.g. via the Hungarian algorithm (Kuhn, 1955). If the resulting assignment doesn’t match the original one, then Knockoffs cannot be exchangeable.
5 Experimental results
We aim at showing the practical consequences of exchangeability violations on error control and diagnosis via C2ST statistics.
Simulated data setup. To assess the consequences of varying degrees of non-exchangeability with Gaussian Knockoffs, we design a simulation setup in which we control the difficulty to produce valid Knockoffs. At each simulation run, we first generate i.i.d. Gaussian data with a tensor structure $\mathbf{X}\in\mathbb{R}^{n\times a\times b\times c}$ with $a\times b\times c=p$. The idea is to generate $n$ samples in a 3D feature space of dimensions $(a,b,c)$. Then, we apply an isotropic smoothing kernel of width $w$ across these three dimensions, mimicking a smooth three-dimensional structure, and flatten the data to obtain $\mathbf{X}\in\mathbb{R}^{n\times p}$.
Then, we draw the true support $\boldsymbol{\beta}^{*}\in\{0,1\}^{p}$. The number of non-null coefficients of $\boldsymbol{\beta}^{*}$ is controlled by the sparsity parameter $s_{p}$, i.e. $s_{p}=\lVert\boldsymbol{\beta}^{*}\rVert_{0}/p$. The target variable $\mathbf{y}$ is built using a linear model:
$\mathbf{y}=\mathbf{X}\boldsymbol{\beta}^{*}+\sigma\boldsymbol{\epsilon},$
where $\sigma=\lVert\mathbf{X}\boldsymbol{\beta}^{*}\rVert_{2}/(\text{SNR}\lVert\boldsymbol{\epsilon}\rVert_{2})$ controls the amplitude of the noise, SNR being the signal-to-noise ratio. We choose the setting $n=500,p=500,s_{p}=0.1,\text{SNR}=2$. To obtain $p=500$, we use $(a,b,c)=(10,10,5)$. Note that using $w=0$ is equivalent to sampling i.i.d. Gaussians since no smoothing is applied in this case. We vary the kernel width $w$ in the interval $[0,1.25]$ to parametrize the difficulty of the problem. This parameter is closely related to the level of correlation observed in the data; as the kernel width increases, the data becomes increasingly correlated.
In all settings, we use the Graphical Lasso for covariance estimation (Friedman, Hastie and Tibshirani, 2008). For each of $N$ simulations, we compute the empirical FDP of the Vanilla Knockoff selection set $S$ and the C2ST statistic averaged across 5-fold cross-validation:
$\widehat{FDP}(S)=\frac{|S\cap\mathcal{H}_{0}|}{|S|},\quad\hat{t}=\frac{1}{n_{\text{te }}}\sum_{\left(z^{i},l^{i}\right)\in\mathcal{D}_{\text{te}}}1\left[g\left(z^{i}\right)=l^{i}\right].$
We first benchmark all existing methods for building Knockoffs for continuous data enumerated in Section 3 using $w=0.5$. This value yields local correlations levels that are comparable to those observed in fMRI data.
Figure 6: C2ST diagnostic metric for varying number of variables and samples. In the left panel, the number of variables $p$ grows and $n=p$ at each point. Note that, when $p>100$ the C2ST accuracy is clearly above chance level for all methods apart from Gaussian Knockoffs and Non-parametric Knockoffs, signaling non-exchangeability in this regime. In the right panel, the number of variables is kept constant with $p$ set to $500$. The number of samples varies from $100$ to $2000$. Apart from Non-parametric Knockoffs, all other methods fail to produce valid Knockoffs, even for large values of $n$.
The left panel of Figure 6 shows that all methods except Gaussian and non-parametric Knockoffs fail when $p$ grows larger (using $n=p$). Amongst the state-of-the-art methods, Gaussian Knockoffs are the best candidate as they maintain a C2ST accuracy of at most $0.6$. The proposed approach maintains chance-level accuracy in all settings. Note that the data at hand is Gaussian – therefore, the non-exchangeability of Gaussian Knockoffs is necessarily due to the problem of covariance estimation. In contrast, the proposed non-parametric approach circumvents this difficult estimation problem by relying on regressions.
The right panel of Figure 6 shows that increasing the number of samples $n$ for a fixed $p$ does not fix the problem, as all methods apart from Non-parametric Knockoffs fail to provide valid Knockoffs. Interestingly, C2ST accuracy rises as $n$ grows larger for all methods apart from Non-parametric Knockoffs. Intuitively, one might expect that as $n$ grows larger, producing valid Knockoffs becomes easier and therefore that the C2ST accuracy should decrease. A possible interpretation of this result is the presence of two competing effects: while learning features of the underlying joint distribution of the data is indeed easier when $n$ grows, the number of training samples accessible to the classifier also grows as it is equal to $2n$. Therefore, the discriminating power of the classifier also improves when $n$ grows larger. From now on, we discard the VAE and Deep approaches because of their poor empirical performance.
Figure 7: FDP and C2ST diagnostic metric for varying smoothing. We use the 3D smoothing kernel width to parametrize the correlation present in the data, which in turn tunes the difficulty to produce valid Knockoffs. For Gaussian Knockoffs, the FDR is controlled only in the easiest settings, i.e. $w\in[0,0.5]$. Note that using $w=0.5$ represents a common setting encountered on real data: it indeed yields local correlations levels that are comparable to those observed in fMRI data. For $w>0.5$ the achieved FDR is substantially above the target FDR, and it grows as the smoothing increases. Note that the C2ST accuracy is clearly above chance level for $w>0.5$, signaling non-exchangeability in this regime. Non-parametric Knockoffs – defined in Section 3.3 – preserve error control in all regimes, which is consistent with C2ST accuracy remaining near chance level.
The left panel of Figure 7 shows that, for Gaussian Knockoffs, the FDR is controlled only in the easiest settings, i.e. $w\in[0,0.5]$. For $w>0.5$ the achieved FDR is substantially above the target FDR, and it grows as the smoothing increases. The $C2ST$ accuracy is clearly above chance level for $w>0.5$. By contrast, non-parametric Knockoffs maintain FDR control in all settings, while the $C2ST$ accuracy remains near chance level.
Semi-simulated data setup. We now turn to evaluating the proposed approach and diagnostic tool on real data. Following Blain et al., 2023; Nguyen, Thirion and Arlot, 2022, we use semi-simulated data to evaluate the proposed method with observed $\mathbf{X}$. We use a simulated response $\mathbf{y}$ to be able to compute the FDP. We consider a first functional Magnetic Resonance Imaging (fMRI) dataset $(\mathbf{X}_{1},\mathbf{y}_{1})$ on which we perform inference using a Lasso estimator; this yields $\boldsymbol{\beta}_{1}^{*}\in\mathbb{R}^{p}$ that we will use as our ground truth. Then, we consider a separate fMRI dataset $(\mathbf{X}_{2},\mathbf{y}_{2})$ for data generation. The point of using a separate dataset is to avoid circularity between the ground truth definition and the inference procedure. Concretely, we discard the original response vector $\mathbf{y}_{2}$ for this dataset and build a simulated response $\mathbf{y}_{2}^{sim}$ using a linear model:
$\mathbf{y}_{2}^{sim}=\mathbf{X}_{2}\boldsymbol{\beta}_{1}^{*}+\sigma\boldsymbol{\epsilon},$
where we set $\sigma$ so that $SNR=4$. We consider $7$ binary classifications problems like "gambling" (rewards vs loss) taken from the HCP dataset. For each of these classification problems, the dataset consists in $778\times 2=1556$ samples and $1000$ features. These features are obtained by averaging fMRI signals within a Ward parcellation scheme, which is known to yield spatially homogeneous regions (Thirion et al., 2014).
As we use $7$ datasets of HCP, we obtain $42=7\times 6$ possible pairs. Since we consider $\boldsymbol{\beta}_{1}^{*}$ as the ground truth, the FDP can be computed. Figure 8 shows False Discovery Proportion levels for 42 semi-simulated fMRI datasets based on HCP data for 5 different Knockoff-based inference methods, using either Gaussian Knockoffs or non-parametric Knockoffs. All methods – see Blain et al., 2023 for a detailed comparison – exhibit problematic False Discoveries proportions using default Gaussian Knockoffs (left panel) and Graphical Lasso-based covariance learning. The expeted behavior is recovered using non-parametric Knockoffs (right panel). As one could expect, the C2ST metric obtains high discrimination power ($0.95$ prediction accuracy) for the Gaussian Knockoffs (signalling non-exchangeability), but is at chance (accuracy of $0.51$) for non-parametric Knockoffs. Regarding the pairing condition, the Hungarian algorithm signals that both Gaussian and Non-parametric Knockoffs are optimally paired with original variables.
Figure 8: Empirical FDP on semi-simulated data for 42 contrast pairs using Gaussian vs Non-parametric Knockoffs. We use 7 HCP contrasts C0: "Motor Hand", C1: "Motor Foot", C2: "Gambling", C3: "Relational", C4: "Emotion", C5: "Social", C6: "Working Memory". We consider all 42 possible train/test pairs: the train contrast is used to obtain a ground truth, while the test contrast is used to generate the response. Inference is performed using the 5 methods considered in the paper and the empirical FDP is reported. Notice that error control of all methods is violated using default Gaussian Knockoffs (left panel) and recovered using Non-parametric Knockoffs (right panel). The C2ST metric is coherent in both cases, with an average accuracy of $0.95$ for the Gaussian Knockoffs (signalling non-exchangeability) and $0.51$ for the non-parametric Knockoffs.
6 Discussion
Knockoffs are a powerful and efficient method for controlled variable selection. This inference procedure allows performing conditional variable selection in one round of inference, without considering each variable individually as in Conditional Randomization Tests approaches.
The statistical guarantees provided by this method rely on the ability to construct valid Knockoffs - namely, Knockoffs and original variables must be exchangeable so that false positives can be demonstrably controlled. However, constructing valid Knockoffs based only on the available observations is not trivial.
The difficulty of this problem is particularly salient in contexts where i) the data at hand has a strong dependence structure and ii) when the number of variables $p$ is large. Note that these are common characteristics of high-dimensional variable selection problems where Knockoffs are relevant, such as fMRI brain mapping or genomic data analyses. These characteristics make the estimation of the joint distribution of the data increasingly difficult, whether done through covariance estimation (as in the Gaussian algorithm) or via deep learning based techniques. Our first contribution in this context is to show that existing generation procedures are unable to produce valid Knockoffs in such settings.
This paper also provides insight into the consequences of exchangeability violations: experiments on both real and simulated data show the important impact of departures from exchangeability on the reliability of Knockoff-based methods. These issues can lead to instability in error control for all Knockoff-based methods. To address this challenge in real-world scenarios, we have introduced a diagnostic tool that provides practitioners with the means to identify exchangeability problems. This tool relies on two parts: a classifier two-sample test which aims at detecting distributional differences between original and Knockoff variables, and a procedure that controls the proper pairing between original observations and their knockoff. Note that having a perfect generative model is not enough to obtain valid Knockoffs, as this does not yield samples paired with the original data. We therefore have two necessary (but not sufficient) conditions, for exchangeability violation.
As part of our efforts to mitigate the exchangeability problems associated with non-Gaussian data or poor covariance estimation, we propose an efficient alternative approach for constructing non-parametric Knockoffs. We prove theoretically that this approach produces valid Knockoffs, provided that the number of samples is large enough. Experiments on simulated data indicate that the Knockoffs we obtain are valid. However, the proposed approach has some limitations. In particular, a sufficient condition for exchangeability is still lacking, so that one may indeed encounter situations where supposedly valid Knockoffs lead to biased FDR guarantees. Another limitation in the non-parametric Knockoff estimation procedure is that the hypothesis that the learner is Bayes optimal leaves open the question of which learner to choose; this choice is obviously important in practice, both for the validity of the Knockoff inference and for the power of the method.
The code for the proposed diagnostic tool and alternative non-parametric Knockoffs algorithm is available at https://github.com/alexblnn/KnockoffsDiagnostics. Note that this algorithm is also less costly in terms of computing time than the Gaussian algorithm. Using a single CPU, a 3×3$\times$3 × speedup is achieved compared to the original algorithm. When leveraging 40 parallel CPUs, a 15×15$\times$15 × speedup is achieved on fMRI data and other large problems. Details about this benchmark are available in Appendix B. Since the proposed algorithm is by design robust to deviations from Gaussianity, it should be preferred in all situations.
Appendix A Proof of Non-parametric Knockoffs validity
Proof of Theorem 1. Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ the original data and $\mathbf{\tilde{X}}\in\mathbb{R}^{n\times p}$ Non-parametric Knockoffs built using Algorithm 1 with a given permutation of $\llbracket p\rrbracket$ denoted $\tau$. Denote $S\subset\llbracket p\rrbracket$ and $j\in\llbracket 2p\rrbracket$. We have, with $(f_{j},\boldsymbol{\widehat{\epsilon}}_{j})$ such that $X_{j}=f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{j}$:
$(\mathbf{X},\mathbf{\tilde{X}})_{j}=\left\{\begin{array}[]{l}f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{j}\text{ if }\quad j\in\llbracket p\rrbracket\\ f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{\tau{(j})}\text{ if }\quad j\in\llbracket p+1,2p\rrbracket\end{array}\right.$
and
$(\mathbf{X},\mathbf{\tilde{X}})_{swap(S),j}=\left\{\begin{array}[]{l}f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{\tau{(j})}\text{ if }\quad j\in S\cap\llbracket p\rrbracket\\ f_{j}(\mathbf{X_{-j}})+\boldsymbol{\widehat{\epsilon}}_{j}\text{ if }\quad j\in S\cap\llbracket p+1,2p\rrbracket\\ (\mathbf{X},\mathbf{\tilde{X}})_{j}\quad\text{ else. }\end{array}\right.$
Let $\mu$ and $\nu$ two distributions of $\mathbb{R}^{2p}$. The usual 2-Wasserstein distance between $\mu$ and $\nu$ is defined as follows, with $\Theta(\mu,\nu)$ the set of distributions of $\mathbb{R}^{2p}\times\mathbb{R}^{2p}$ with marginals $\mu$ and $\nu$:
$W_{2}(\mu,\nu):=\left(\inf_{\theta\in\Theta(\mu,\nu)}\int_{\mathbb{R}^{2p}\times\mathbb{R}^{2p}}||x-y||^{2}\mathrm{d}\theta(x,y)\right)^{1/2}$
It follows that $W_{2}((\mathbf{X},\mathbf{\tilde{X}})_{swap(S)},(\mathbf{X},\mathbf{\tilde{X}}))\leq W_{2}(\widehat{\boldsymbol{\epsilon}},\widehat{\boldsymbol{\epsilon}}_{\tau})\leq W_{2}(\widehat{\boldsymbol{\epsilon}},\boldsymbol{\epsilon})+W_{2}(\widehat{\boldsymbol{\epsilon}}_{\tau},\boldsymbol{\epsilon}_{\tau})$ with $\mathbf{\epsilon}$ the true noise vector such that:
$\forall j\in\llbracket p\rrbracket,\quad X_{j}=f^{*}_{j}(\mathbf{X_{-j}})+\boldsymbol{\epsilon}_{j}.$
[…]
$W_{2}((\mathbf{X},\mathbf{\tilde{X}})_{swap(S)},(\mathbf{X},\mathbf{\tilde{X}}))\xrightarrow[\enskip n\to\infty\enskip]{}0$
Since $W_{2}$ is a distance over distributions in $\mathbb{R}^{2p}$, this proves that Non-parametric Knockoffs are exchangeable asymptotically in the number of samples $n$.
Appendix B Computation time of non-parametric Knockoffs
We have performed benchmarks to evaluate the computation time needed for the proposed Non-parametric Knockoffs algorithm compared to the Gaussian algorithm. We use the simulation setup described in Section 4 and vary the number of variables $p$. For all values of $p$, we use $n=p$ samples. We first consider using a single CPU and then consider using 40 CPUs. For the Gaussian approach, parallelization is done in the Graphical Lasso covariance estimator via the $n_{jobs}$ argument in scikit-learn (Pedregosa et al., 2011). Note that profiling the code when using a single CPU shows that most of the computation time is spent in covariance estimation. For the Non-parametric approach, the 40 CPUs are used to train Lasso models that predict $X_{j}$ from $\mathbf{X_{-j}}$ in parallel.
Figure 9: Non-parametric and Gaussian Knockoffs computation time. We use either one CPU or 40 CPUs and report the computation time for various problem dimensions in log-log scale. Note that using non-parametric Knockoffs on a single CPU yields around a 3×3$\times$3 × speedup across all problem dimensions compared to Gaussian Knockoffs. Using 40 CPUs, a 5×5$\times$5 × speedup is achieved for problems under $1000$ variables, and a 10×10$\times$10 × to 15×15$\times$15 × speedup for problems of up to $2000$ variables.
Note that using non-parametric Knockoffs on a single CPU yields around a 3×3$\times$3 × speedup across all problem dimensions compared to Gaussian Knockoffs. Using 40 CPUs, a 5×5$\times$5 × speedup is achieved for problems under $1000$ variables, and a 10×10$\times$10 × to 15×15$\times$15 × speedup for problems of up to $2000$ variables.